# Probing metric fluctuations with the spin of a particle: a quantum simulation with bimodal optical cavities

### v1

@@ Abstract
Exploring potential empirical manifestations of quantum gravity is a challenging pursuit. In this study, we utilise a lattice representation of a $(2+1)$D massive gravity toy model interacting with Dirac fermions that can support specific spacetime fluctuations. We focus on the evolution of the fermion’s spin due to its coupling to spacetime fluctuations. To monitor their dynamics a minimal model is required that comprises two bosonic modes describing spacetime geometry fluctuations coupled with the spin of the fermion. A possible emulation of this system involves encoding spin degrees of freedom in an atom coupled with a bimodal optical cavity that provides the two bosonic modes. We observe diverse spin evolution patterns based on its interaction with fluctuating spacetime geometry. Our proposal introduces a novel approach for modelling the effect of interactions between quantum gravity and matter that can be probed with current technology.
@@ 1 Introduction
Another significant area of exploration involves certain collective excitations in fractional quantum Hall liquids, which remarkably behave as analogue gravitons. Specifically, the Girvin-MacDonald-Platzman mode of the fractional quantum Hall effect [5, 6], often termed the magnetoroton, is recognized as a massive graviton excitation in its long-wavelength limit [7, 8, 9]. Experimental observations have corroborated these excitations [10, 11, 12]. However, the strong interactions inherent to fractional quantum Hall liquids make them analytically intractable, highlighting the need for alternative approaches that can emulate their emergent features.
@@
To overcome these analytical limitations, an alternative strategy involves engineering systems capable of effectively simulating quantum signatures of gravity. Such quantum simulations offer the advantage of accessing a broad spectrum of coupling regimes, potentially amplifying properties that are otherwise experimentally inaccessible. Recent progress in simulating quantum gauge theories, for instance, has provided valuable insights into complex phenomena like quark confinement [13]. While various platforms exist for simulating classical gravity, such as deforming graphene-like systems to produce non-trivial extrinsic geometry [14, 15] or creating tunneling coupling inhomogeneities that generate intrinsic geometries [16, 17], simulating quantum gravity has remained an open challenge. Roadblocks persist at both conceptual levels, like realizing quantum fluctuations of spacetime, and practical levels, such as engineering the intricate self-interactions characteristic of gravitational theories.
@@
This work is motivated by the importance of bridging the gap between theory and experiment in quantum gravity. We propose an experimental setup that can simulate quantum metric fluctuations, thus facilitating the identification of novel physical manifestations of quantum gravity. Our starting point is based on a lattice representation of a particular model for metric perturbations in $(2+1)$ dimensions coupled to Dirac fermions [18]. Unlike the model considered in [18], the fluctuations of the spatial geometry considered here are described by a traceless symmetric tensor, which can be associated to a unimodular spatial metric. Therefore, our lattice model gives rise to non-relativistic massive metric perturbations similar to the massive gravitons identified in the study of the Girvin–MacDonald–Platzman (GMP) mode in the fractional quantum Hall effect. [8, 9]. This framework allows us to focus on a simplified set of spacetime fluctuations that still exhibit a rich dynamical behaviour. We employ the interactions between matter and spacetime fluctuations to use the spin of the Dirac fermions as a probe to deduce the behaviour of the gravitational dynamics.
@@
The lattice representation of quantum gravity coupled to matter has also the advantage of being readily amenable to experimental simulations, thus bridging the gap between theory and experiment in this field of research. In particular, we identify the smallest instance of the lattice model that encodes the coupling between spacetime fluctuations and the spin of the Dirac fermion. We then propose an experimental setup for the quantum simulation of this minimal model. This setup involves encoding the spin degrees of freedom in the atomic levels of a single atom, placed inside a bimodal optical cavity that encodes two bosonic modes responsible for the spacetime fluctuations. By varying the coupling between the atom and the cavity we are able to use the atom as a probe to identify the dynamics of the simulated spacetime fluctuations. We identify distinct dynamical signatures of the spin across different gravitational regimes, illustrating how spacetime fluctuations influence matter evolution at the quantum level. These findings open a pathway for experimental probes of quantum gravitational phenomena possibly by enhancing the effect fluctuating gravity can have on matter. Moreover, our research establishes a platform for simulating more intricate models of gravity–matter interactions in future work. Beyond their fundamental significance, such quantum simulations could shed light on information flow, entanglement dynamics, and decoherence in curved quantum geometries, offering valuable insights for both quantum gravity and quantum technologies.
@@ 2 Lattice representation of Dirac field coupled to a massive spin-2 field
We now consider a massless Dirac fermion field coupled to metric perturbations described by a toy model Hamiltonian for a bosonic massive spin-2 field. Subsequently, we will consider the lattice model of this composite system.
@@ 2.1 Carrollian massive spin-2 field
Moreover, we restrict $h_{ij}$ to be traceless i.e., $h\equiv\delta^{ij}h_{ij}=0$. This has the consequence that the metric determinant is fixed to be ${\rm det}(g)=1$, a characteristic of unimodular gravity models. Since Einstein gravity in 2+1 dimensions lacks propagating degrees of freedom [19], the massless Fierz-Pauli action coming from linearising the Einstein-Hilbert Lagrangian has trivial dynamics for $h_{ij}$. To introduce non-trivial dynamics, we instead model $h_{ij}$ as a massive spin-2 field. In contrast, the adopted Fierz–Pauli action for a relativistic massive spin-2 field in 2+1 dimensions, describes two local propagating degrees of freedom [20], thus having a very similar behaviour to the linearised Einstein gravity in four dimensions. For our purposes, it is sufficient to focus on the temporal dynamics of the metric fluctuations, neglecting spatial derivatives of $h_{ij}$. We therefore consider the following action principle
[…]
@@
To quantise the model, $h_{ij}$ and $p_{ij}$ are promoted to operators with commutation relations compatible with the Dirac bracket {,}$\{\,,\,\}^{*}\rightarrow-i[\,,\,]$. Equivalently, we parametrise the independent components of $h_{ij}$ and $p_{ij}$ in terms of operators $\tilde{\alpha}$ and $\tilde{\beta}$ satisfying the commutation relations
$[\tilde{\alpha}(\boldsymbol{x}),\tilde{\alpha}^{\dagger}(\boldsymbol{x}^{\prime})]=\delta^{(2)}(\boldsymbol{x}-\boldsymbol{x}^{\prime})=[\tilde{\beta}(\boldsymbol{x}),\tilde{\beta}^{\dagger}(\boldsymbol{x}^{\prime})],$ (8)
@@
$\tilde{\alpha}(\boldsymbol{k})=\alpha(\boldsymbol{k})\cosh{r}-\alpha^{\dagger}(-\boldsymbol{k})\sinh r,\,\,\,\,\,\,\,\,\tilde{\beta}(\boldsymbol{k})=\beta(\boldsymbol{k})\cosh r-\beta^{\dagger}(-\boldsymbol{k})\sinh r,$ (11)
@@
$\cosh{2r}=\frac{\mu}{4}+\frac{1}{\mu},\qquad\sinh{2r}=\frac{\mu}{4}-\frac{1}{\mu},$ (12)
@@ 2.2 Coupling to Dirac fermion
describing $(2+1)$D Dirac fermions coupled to semiclassical gravity fluctuations. In the following we will present a lattice model that gives rise to this Hamiltonian in the continuum limit.
@@ 2.3 Lattice Hamiltonian
Figure 1: (Left) The brick wall lattice representation of the $(2+1)$D Dirac fermions coupled to spacetime fluctuations given by Hamiltonian (18). The unit cell includes two fermionic modes, $a$ and $b$, that encode the spin of the Dirac field. The bosonic modes $\tilde{\alpha}$ and $\tilde{\beta}$ describe spacetime fluctuations and are coupled to the fermions through their tunnelling. The basis vectors $\boldsymbol{n}_{1}$ and $\boldsymbol{n}_{2}$ span the brick wall lattice of fermions and $J_{x}$, $J_{y}$ and $J_{z}$ are the fermionic tunnelling couplings that depend on the bosonic modes $\tilde{\alpha}$ and $\tilde{\beta}$. (Right) The minimal system considered here with two fermionic modes that encode a single spin and two bosonic modes coupled to the fermions.
The $(2+1)$D Dirac fermion coupled to the massive spin-2 field can effectively arise from a 2D lattice model. In [18] a two-dimensional lattice of interacting bosonic and fermionic modes is presented that gives rise to a Hamiltonian that resembles (16) in its low-energy limit. A similar strategy is adopted here where we start with a brick wall lattice of fermions, as shown in Fig. 1 (Left). The lattice Hamiltonian $H_{\rm latt}$ is given by
@@
$[a^{\dagger}_{\boldsymbol{i}},a_{\boldsymbol{j}}]=2\delta_{\boldsymbol{i},\boldsymbol{j}}=[b^{\dagger}_{\boldsymbol{i}},b_{\boldsymbol{j}}],$ (21)
@@
$H_{\rm f+b}=\sum_{\boldsymbol{i}}\bigg{(}a^{\dagger}_{\boldsymbol{i}}J_{X}(\boldsymbol{i})b_{\boldsymbol{i}+\boldsymbol{n}_{1}}+a^{\dagger}_{\boldsymbol{i}}J_{Y}(\boldsymbol{i})b_{\boldsymbol{i}+\boldsymbol{n}_{2}}+a^{\dagger}_{\boldsymbol{i}}J_{Z}(\boldsymbol{i})b_{\boldsymbol{i}}\bigg{)}+{\rm h.c.},$ (22)
@@
Following a similar procedure as in [18] one can show that the lattice of tunnelling fermions and bosons shown in Fig. 1 (Left) gives rise, in its low energy limit, to a $(2+1)$D Dirac fermion coupled to massive metric fluctuations, as given in (16). While it is straightforward to see the continuum limit of (20) in momentum space leads to Eq. (10), the fermionic Hamiltonian requires some extra analysis. First, one should notice that the Fermi points of the free Hamiltonian ($J_{X}=J_{Y}=J_{Z}/\sqrt{2}=1$) read
@@
$H_{\rm f+b}=\sum_{\boldsymbol{p},\boldsymbol{q}}\Psi^{+\dagger}_{\boldsymbol{p}}\,h^{+}(\boldsymbol{p},\boldsymbol{q})\,\Psi^{+}_{\boldsymbol{q}}+\sum_{\boldsymbol{p},\boldsymbol{q}}\Psi^{-\dagger}_{\boldsymbol{p}}\,h^{-}(\boldsymbol{p},\boldsymbol{q})\,\Psi^{-}_{\boldsymbol{q}},$ (26)
@@
allows one to write $H_{\rm f+b}$ in Dirac form $\int d^{2}x\;\psi^{\dagger}h\,\psi$. When expressing the modes $\tilde{\alpha}$ and $\tilde{\beta}$ in terms of the metric perturbations as given in Eq. (9), the single-particle Hamiltonian $h$ matches precisely (17).
3 Spin-$1/2$ coupled to local geometry fluctuations
In the previous section we presented the lattice Hamiltonian that effectively encodes the dynamics of a Dirac field coupled to fluctuating geometric background. A proposal of how to realise a similar lattice model with optical lattices and a mixture of bosonic and fermionic atoms was given in [18]. Here we want to consider only the evolution of the Dirac fermion’s spin due to the fluctuating background. The minimal subsystem of the lattice that is required to encode and monitor this evolution comprises two fermionic modes that encode the spin of the particle and the two bosonic eigenmodes that couple resonantly to the fermionic ones. Such a minimal system provides an analytically tractable way to predict the time evolution of the system. Moreover, it lends itself for quantum simulations with optical cavity technology.
3.1 Minimal model
We can isolate the dynamics that the spin-2 field induces on the spin-$1/2$ of the Dirac particle by considering a single fermionic unit cell with $\boldsymbol{i}={\mathbf{0}}$. In this case the fermion-boson Hamiltonian reduces to
$H_{\text{f}+\text{b}}=J_{Z}a^{\dagger}_{\mathbf{0}}b_{\mathbf{0}}+J_{Z}^{\dagger}b^{\dagger}_{\mathbf{0}}a_{\mathbf{0}},$ (30)
@@ 3.1 Minimal model
$J_{Z}=\sqrt{2}+2i\sqrt{\pi G}(\tilde{\alpha}_{\mathbf{0}}+\tilde{\alpha}^{\dagger}_{\mathbf{0}})-2\sqrt{\pi G}(\tilde{\beta}_{\mathbf{0}}+\tilde{\beta}^{\dagger}_{\mathbf{0}}).$ (31)
@@
$\tilde{\alpha}_{\mathbf{0}}=\frac{1}{\sqrt{V}}\sum_{\boldsymbol{k}}e^{i\boldsymbol{k}\cdot\boldsymbol{i}}\tilde{\alpha}_{\boldsymbol{k}}\bigg{|}_{\boldsymbol{i}=\mathbf{0}}\rightarrow\int\frac{d^{2}k}{2\pi}\tilde{\alpha}(\boldsymbol{k})$ (32)

### v2

# Probing metric fluctuations with the spin decoherence of a particle in a quantum simulation

@@ Abstract
Exploring potential empirical manifestations of quantum gravity is a challenging pursuit. In this study, we utilise a lattice representation of a (2+1)(2+1)D massive gravity toy model interacting with Dirac fermions that can support specific spacetime fluctuations. We focus on the evolution of the fermion’s spin due to its coupling to spacetime fluctuations. To monitor their dynamics a minimal model is required that comprises two bosonic modes describing spacetime geometry fluctuations coupled with the spin of the fermion. A possible emulation of this system involves encoding spin degrees of freedom in the electronic states of an atom coupled with a bimodal optical cavity that provides the two bosonic modes. The spin exhibits a variety of dynamical behaviours due to its coupling with the fluctuating geometry, with decoherence emerging as a key signature of this interaction. Our proposal introduces a novel approach for modelling the effect of interactions between quantum gravity and matter that can be probed with current technology.
@@ 1 Introduction
Another significant area of exploration involves certain collective excitations in fractional quantum Hall liquids, which remarkably behave as analogue gravitons. Specifically, the Girvin-MacDonald-Platzman mode of the fractional quantum Hall effect [5, 6], often termed the magnetoroton, is recognized as a massive graviton excitation in its long-wavelength limit [7, 8, 9]. Moreover, the predicted properties of these excitations are in agreement with current experimental observations [10, 11, 12]. However, the strong interactions inherent to fractional quantum Hall liquids make them analytically intractable, highlighting the need for alternative approaches that can emulate their emergent features.
@@
To overcome these analytical limitations, an alternative strategy involves engineering systems capable of effectively simulating quantum signatures of gravity. Such quantum simulations offer the advantage of accessing a broad spectrum of coupling regimes. In particular, they enable to amplify properties of gravity that are otherwise experimentally inaccessible due to its very weak coupling to matter. Recent progress in simulating quantum gauge theories, for instance, has provided valuable insights into complex phenomena like quark confinement [13]. While various platforms exist for simulating classical gravity, such as deforming graphene-like systems to produce non-trivial extrinsic geometry [14, 15] or creating tunneling coupling inhomogeneities that generate intrinsic geometries [16, 17], simulating quantum gravity has remained an open challenge. Roadblocks persist at both conceptual levels, like realizing quantum fluctuations of spacetime, and practical levels, such as engineering the intricate self-interactions characteristic of gravitational theories.
@@
This work is motivated by the importance of bridging the gap between theory and experiment in quantum gravity. We propose an experimental setup that can simulate quantum metric fluctuations, thus facilitating the identification of novel physical manifestations of quantum gravity. Our starting point is based on a lattice representation of a particular model for metric perturbations in $(2+1)$ dimensions coupled to Dirac fermions [18]. Unlike the model considered in [18], the fluctuations of the spatial geometry considered here are described by a traceless symmetric tensor, which can be associated to a unimodular spatial metric. Therefore, our lattice model gives rise to non-relativistic massive metric perturbations similar to the massive gravitons identified in the study of the Girvin–MacDonald–Platzman (GMP) mode in the fractional quantum Hall effect [8, 9]. This framework allows us to focus on a simplified set of spacetime fluctuations that still exhibit a rich dynamical behaviour. We employ the interactions between matter and spacetime fluctuations to use the spin of the Dirac fermions as a probe to deduce the behaviour of the gravitational dynamics.
@@
The lattice representation of quantum gravity coupled to matter has also the advantage of being readily amenable to experimental simulations, thus bridging the gap between theory and experiment in this field of research. In particular, we identify the smallest instance of the lattice model that encodes the coupling between spacetime fluctuations and the spin of the Dirac fermion. We then propose an experimental setup for the quantum simulation of this minimal model. This setup involves encoding the spin degrees of freedom in the electronic levels of a single atom, placed inside a bimodal optical cavity that encodes two bosonic modes responsible for the spacetime fluctuations. By varying the coupling between the atom and the cavity we are able to use the atom as a probe to identify the dynamics of the simulated spacetime fluctuations. We identify distinct dynamical signatures of the spin across different gravitational regimes, illustrating how spacetime fluctuations influence matter evolution at the quantum level. These findings open a pathway for experimental probes of quantum gravitational phenomena possibly by enhancing the effect fluctuating gravity can have on matter. Moreover, our research establishes a platform for simulating more intricate models of gravity–matter interactions in future work. Beyond their fundamental significance, such quantum simulations could shed light on information flow, entanglement dynamics, and decoherence in curved quantum geometries, offering valuable insights for both quantum gravity and quantum technologies.
@@ 2 Lattice representation of Dirac field coupled to a massive spin-2 field
We now consider a massless Dirac fermion field coupled to metric perturbations described by a toy model Hamiltonian for a bosonic massive spin-2 field. Subsequently, we will consider a lattice model that describes this composite system in its low energy limit.
@@ 2.1 Carrollian massive spin-2 field
Moreover, we restrict $h_{ij}$ to be traceless i.e., $h\equiv\delta^{ij}h_{ij}=0$. This has the consequence that the metric determinant is fixed to be ${\rm det}(g)=1$, a characteristic of unimodular gravity models. Since Einstein gravity in 2+1 dimensions lacks propagating degrees of freedom [19], the massless Fierz-Pauli action coming from linearising the Einstein-Hilbert Lagrangian has trivial dynamics for $h_{ij}$. To introduce non-trivial dynamics, we instead model $h_{ij}$ as a massive spin-2 field. The adopted Fierz–Pauli action for a relativistic massive spin-2 field in 2+1 dimensions, describes two local propagating degrees of freedom [20], thus having a very similar behaviour to the linearised Einstein gravity in four dimensions. For our purposes, it is sufficient to focus on the temporal dynamics of the metric fluctuations, neglecting spatial derivatives of $h_{ij}$, while still capturing the essential physical behaviour and complexity of the system. We therefore consider the following action principle
$S_{\rm gr}=\frac{1}{64\pi G}\int dt\int d^{2}x\;\bigg(-\dot{h}_{ij}\dot{h}^{ij}-\dot{h}^{2}+\mu^{2}\left(h_{ij}h^{ij}-h^{2}\right)+\xi\,h\bigg),$ (2)
@@
To quantise the model, $h_{ij}$ and $p_{ij}$ are promoted to operators with commutation relations compatible with the Dirac bracket $\{\,,\,\}^{*}\rightarrow-i[\,$,\,]. Equivalently, we parametrise the independent components of $h_{ij}$ and $p_{ij}$ in terms of operators $\tilde{\alpha}$ and $\tilde{\beta}$ satisfying the commutation relations
$\tilde{\alpha}(\boldsymbol{x}),\tilde{\alpha}^{\dagger}(\boldsymbol{x}^{\prime})]=\delta^{(2)}(\boldsymbol{x}-\boldsymbol{x}^{\prime})=[\tilde{\beta}(\boldsymbol{x}),\tilde{\beta}^{\dagger}(\boldsymbol{x}^{\prime})],$ (8)
@@
[…]
@@
[…]
@@ 2.2 Coupling to Dirac fermion
describing (2+1)(2+1)D Dirac fermions coupled to semiclassical gravity fluctuations. In the following we will present a lattice model that gives rise to this Hamiltonian in the continuum limit.
@@ 2.3 Lattice Hamiltonian
Figure 1: (Left) The brick wall lattice representation of the (2+1)(2+1)D Dirac fermions coupled to spacetime fluctuations given by Hamiltonian (18). The unit cell includes two fermionic modes, $a$ and $b$, that encode the spin of the Dirac field. The bosonic modes $\tilde{\alpha}$ and $\tilde{\beta}$ describe spacetime fluctuations and are coupled to the fermions through their tunnelling. The basis vectors $\boldsymbol{n}_{1}$ and $\boldsymbol{n}_{2}$ span the brick wall lattice of fermions and $J_{x}$, $J_{y}$ and $J_{z}$ are the fermionic tunnelling couplings that depend on the bosonic modes $\tilde{\alpha}$ and $\tilde{\beta}$. (Right) The minimal system considered here with two fermionic modes that encode a single spin and two bosonic modes coupled to the fermions.
The (2+1)(2+1)D Dirac fermion coupled to the massive spin-2 field can effectively arise from a 2D lattice model. In [18] a two-dimensional lattice of interacting bosonic and fermionic modes is presented that gives rise to a Hamiltonian that resembles (16) in its low-energy limit. A similar strategy is adopted here where we start with a brick wall lattice of fermions, as shown in Fig. 1 (Left). The lattice Hamiltonian $H_{\rm latt}$ is given by
@@
$a^{\dagger}_{\boldsymbol{i}},a_{\boldsymbol{j}}]=2\delta_{\boldsymbol{i},\boldsymbol{j}}=[b^{\dagger}_{\boldsymbol{i}},b_{\boldsymbol{j}}],$ (21)
@@
$H_{\text{f-b}}=\sum_{\boldsymbol{i}}\bigg(a^{\dagger}_{\boldsymbol{i}}J_{X}(\boldsymbol{i})b_{\boldsymbol{i}+\boldsymbol{n}_{1}}+a^{\dagger}_{\boldsymbol{i}}J_{Y}(\boldsymbol{i})b_{\boldsymbol{i}+\boldsymbol{n}_{2}}+a^{\dagger}_{\boldsymbol{i}}J_{Z}(\boldsymbol{i})b_{\boldsymbol{i}}\bigg)+{\rm h.c.},$ (22)
@@
Following a similar procedure as in [18] one can show that the lattice of tunnelling fermions and bosons shown in Fig. 1 (Left) gives rise, in its low energy limit, to a (2+1)(2+1)D Dirac fermion coupled to massive metric fluctuations, as given in (16). While it is straightforward to see the continuum limit of (20) in momentum space leads to Eq. (10), the fermionic Hamiltonian requires some extra analysis. First, one should notice that the Fermi points of the free Hamiltonian ($J_{X}=J_{Y}=J_{Z}/\sqrt{2}=1$) read
@@
$H_{\text{f-b}}=\sum_{\boldsymbol{p},\boldsymbol{q}}\Psi^{+\dagger}_{\boldsymbol{p}}\,h^{+}(\boldsymbol{p},\boldsymbol{q})\,\Psi^{+}_{\boldsymbol{q}}+\sum_{\boldsymbol{p},\boldsymbol{q}}\Psi^{-\dagger}_{\boldsymbol{p}}\,h^{-}(\boldsymbol{p},\boldsymbol{q})\,\Psi^{-}_{\boldsymbol{q}},$ (26)
@@
allows one to write $H_{\text{f-b}}$ in Dirac form $\int d^{2}x\;\psi^{\dagger}h\,\psi$. When expressing the modes $\tilde{\alpha}$ and $\tilde{\beta}$ in terms of the metric perturbations as given in Eq. (9), the single-particle Hamiltonian $h$ matches precisely (17).
3 Dynamical evolution of a spin-$1/2$ due to geometry fluctuations
In the previous section we presented the lattice Hamiltonian that effectively encodes the dynamics of a Dirac field coupled to fluctuating geometric background. A proposal of how to realise a similar lattice model with optical lattices and a mixture of bosonic and fermionic atoms was given in [18]. Here we want to consider only the evolution of the Dirac fermion’s spin due to the fluctuating background. The minimal subsystem of the lattice that is required to encode and monitor this evolution comprises two fermionic modes that encode the spin of the particle and the two bosonic eigenmodes that couple resonantly to the fermionic ones. Such a minimal system provides an analytically tractable way to predict the time evolution of the system. Moreover, it lends itself for quantum simulations with optical cavity technology as we shall see in this section.
3.1 Spin-$1/2$ coupled to quantum geometry fluctuations
3.1.1 Minimal model
We can isolate the dynamics that the spin-2 field induces on the spin-$1/2$ of the Dirac particle by considering a single fermionic unit cell with $\boldsymbol{i}={\mathbf{0}}$. Omitting the site label, the fermion-boson Hamiltonian reduces to
$H_{\text{f-b}}=J_{Z}a^{\dagger}b+J_{Z}^{\dagger}b^{\dagger}a,$ (30)
@@ 3.1.1 Minimal model
[…]
@@
$\tilde{\alpha}=\int\frac{d^{2}k}{2\pi}\tilde{\alpha}(\boldsymbol{k})$ (32)
@@
$H=\sqrt{2}a^{\dagger}b+\int_{0}^{\infty}kdk\Bigg[2\pi\mu\left(\alpha^{\dagger}_{0}\alpha_{0}+\beta^{\dagger}_{0}\beta_{0}\right)-2\sqrt{\frac{\pi G}{\mu}}\big(\beta_{0}+\beta^{\dagger}_{0}-i\alpha_{0}-i\alpha^{\dagger}_{0}\big)a^{\dagger}b\Bigg]+{\rm h.c.},$ (34)
where we have neglected the Fourier modes $\alpha_{m}(k)$ and $\beta_{m}(k)$ for $m\neq 0$ since they decouple from the fermonic modes. Out of all the possible momenta, the dominant terms in this Hamiltonian are the ones for which the resonant condition $\sqrt{2}\approx 2\pi k\mu$ holds, which defines a circle in momentum space of radius $k_{R}=\frac{1}{\sqrt{2}\pi\mu}$. Dropping the operator labels for the bosonic modes, the Hamiltonian can be rewritten as
$H=\sqrt{2}\sigma^{x}+\sqrt{2}\left(\alpha^{\dagger}\alpha+\beta^{\dagger}\beta\right)+g_{\text{s-b}}\left[\left(\beta+\beta^{\dagger}\right)\sigma^{x}+\left(\alpha+\alpha^{\dagger}\right)\sigma^{y}\right],$ (35)
$g_{\text{s-b}}=-\frac{\sqrt{2G}}{\sqrt{\pi}\mu^{3/2}}.$ (36)
3.2 Dynamical evolution of spin